### v1

# Altermagnetic Spin Precession and Spin Transistor

Abstract
Altermagnets hold great potential for spintronic applications, yet their intrinsic spin dynamics and associated transport properties remain largely unexplored. Here, we investigate spin-resolved quantum transport in a multi-terminal setup based on a $d$-wave altermagnet. It is found that the altermagnetic spin splitting in momentum space induces an interesting spin precession in real space, giving rise to characteristic spin patterns. This altermagnetic spin precession manifests as a spatial modulation of the Hall voltage, whose oscillation period provides a direct measure of the spin-splitting strength. When the altermagnetism is electrically tunable, the proposed setup functions as a prototype for a highly efficient spin transistor. The key physical effects are shown to be robust against dephasing and crystalline warping. Our work not only identifies a fingerprint signature of altermagnets, offering a direct probe of the altermagnetic spin splitting, but also represents an important step toward bridging their fundamental physics with practical spintronic applications.
Introduction.—Various magnetic phases are conventionally classified based on the arrangement of magnetic atoms, with the role of nonmagnetic constituents typically neglected neel1971magnetism; vsmejkal2022emerging. However, recent studies on altermagnetism have revealed that nonmagnetic atoms can play a significant role in shaping the magnetic properties of the system vsmejkal2022emerging. In particular, the antiferromagnetic sublattices with opposite spins cannot be related by inversion or translation operations vsmejkal2022beyond, giving rise to pronounced momentum-dependent spin splitting in the band structure, despite the absence of net magnetization wu2007fermi; hayami2020bottom; naka2021perovskite; yuan2021prediction; yuan2020giant; naka2019spin; hayami2019momentum; ahn2019antiferromagnetism; vsmejkal2022anomalous; ma2021multifunctional; mazin2021prediction. Notably, the spin splitting in altermagnets can exceed that induced by spin-orbit coupling by an order of magnitude and exhibits unconventional $d$-, $g$-, and $i$-wave symmetries vsmejkal2022beyond; krempasky2024altermagnetic; lin2024observation, opening up promising opportunities for novel spintronic applications.
Figure 1: (a) Schematic of the proposed setup, consisting of a central altermagnetic region connected to two longitudinal leads (1 and 2) and multiple transverse voltage probes. Top and bottom gates ($V_{tg}$ and $V_{bg}$) control the altermagnetic splitting, which is essential for the spin transistor. The $d$-wave altermagnetic Fermi surfaces are illustrated in the inset. (b) Altermagnetic spin precession in real space, with arrows indicating the local spin orientations. (c,d) Spatial distribution of spin component $S_{y}$ for nonequilibrium propagating electrons under (c) point injection and (d) line injection, with the injected spins polarized along the $x$ direction. The solid and dashed arrows highlight the corresponding spin modulation along the $x$ direction in (c) and (d). A square-root scale is applied to the colorbar to enhance the contrast around zero.
Significant progress has been made in the search for altermagnetic materials, both theoretically and experimentally bai2024altermagnetism; song2025altermagnets. In particular, first-principles calculations have predicted hundreds of potential altermagnetic candidates bai2024altermagnetism; sodequist2024two; vsmejkal2022beyond; vsmejkal2022emerging; jiang2025metallic; zhang2025crystal; hu2025catalog; chen2024enumeration, and altermagnetic spin splitting has been experimentally observed through angle-resolved photoemission spectroscopy zhu2024observation; fedchenko2024observation; krempasky2024altermagnetic; ding2024large; lin2024observation; reimers2024direct; jiang2025metallic; zhang2025crystal; yang2025three; jungwirth2025symmetry. A more refined characterization of the altermagnetic Fermi surface, through quantum oscillation measurements li2025diagnosing and quasiparticle interference imaging hu2025quasiparticle; lou24prb; linder24prb, is warranted. The distinctive band structures of altermagnets can lead to a range of notable physical phenomena, including the anomalous Hall effect vsmejkal2020crystal; feng2022anomalous; vsmejkal2022anomalous; reichlova2024observation; gonzalez2023spontaneous; zhou2025manipulation; galindez2025revealing; leiviska2024anisotropy; li2025exploration; yu2025neel, giant magnetoresistance vsmejkal2022beyond; vsmejkal2022giant; li2025marginal, magnetic nonlinear Hall effect han2025discovery, spin-charge conversion karube2022observation; bai2023efficient; gonzalez2021efficient; jechumtal2025spin, interesting Josephson effects li2025spin; zhang2024finite, novel magnetoelectric coupling chen2025electrical; duan2025antiferroelectric; gu2025ferroelectric and Hall drag effects lin2025coulomb. While these advances highlight the great potential of altermagnets for spintronic applications, the electron spin dynamics and the resulting quantum transport properties remain poorly understood.
In this work, we address this gap by investigating spin-resolved quantum transport in a multi-terminal device based on a two-dimensional $d$-wave altermagnet, as illustrated in Fig. 1(a). Owing to the altermagnetic spin splitting, electrons injected with spin polarization perpendicular to the Néel vector of the altermagnet undergo a characteristic spin precession [Fig. 1(b)], resulting in distinctive real-space spin patterns [Figs. 1(c,d)]. These spin patterns faithfully reflect the underlying altermagnetic spin splitting in momentum space and can be probed through the spatial oscillations of the Hall voltage. The oscillation period, in particular, is an analytical function of the altermagnetic spin splitting, providing a direct measure of this key physical parameter. The physical effects are robust against both dephasing and crystalline warping. Moreover, for electrically tunable altermagnets, the proposed setup naturally functions as a prototype for a high-efficiency spin transistor. Our results establish a fingerprint signature of altermagnets and pave the way for their device applications.
Altermagnetic spin precession.—We consider a two-dimensional $d$-wave altermagnet, specifically with $d_{xy}$-type order, to illustrate the unique spin precession mechanism. The system is modeled by the effective Hamiltonian as vsmejkal2022beyond
$H=\frac{B}{2}\left(k_{x}^{2}+k_{y}^{2}\right)+\alpha_{A}k_{x}k_{y}\sigma_{z}-\mu,$ (1)
where $B$ denotes the inverse effective mass of the electron, $\alpha_{A}$ characterizes the altermagnetic spin-splitting, $\sigma_{z}$ (and later $\sigma_{x,y}$) are the Pauli matrices representing electron spin, and $\mu$ is the chemical potential.
Similar to the spin-orbit coupling effect, the altermagnetic spin splitting can be viewed as a momentum-resolved Zeeman field, leading to distinctive spin precession and, consequently, characteristic spin patterns in real space. To illustrate this, we first consider the local injection of electrons at the origin without loss of generality. The propagator from the origin to $\bm{r}$ takes the form SM
$g^{R}_{\sigma,\sigma^{\prime}}(\bm{r};\omega)=-\frac{4\pi\delta_{\sigma,\sigma^{\prime}}}{\sqrt{B^{2}-\alpha_{A}^{2}}}K_{0}(-i{\bm{k}}_{\sigma}\cdot\bm{r}),$ (2)
where the subscripts $\sigma,\sigma^{\prime}=\uparrow,\downarrow$ label the spin components and $K_{0}(\cdot)$ is the zeroth-order modified Bessel function of the second kind. The Kronecker delta $\delta_{\sigma,\sigma^{\prime}}$ indicates the absence of spin flipping during electron propagation, a consequence of the collinear spin splitting in altermagnets. For a given energy $\omega$, the wavevector is spin-dependent and can be parameterized as $\bm{k}_{\sigma}(\omega,\theta)=k_{x,\sigma}(1,f_{\sigma})$ where $k_{x,\sigma}(\theta)=\sqrt{(\omega+\mu)/[B(1+f_{\sigma}^{2})/2+\sigma\alpha_{A}f_{\sigma}]}$ and $f_{\sigma}(\theta)=-f_{-\sigma}(-\theta)=(B\tan\theta-\sigma\alpha_{A})/(B-\sigma\alpha_{A}\tan\theta)$, and the propagation angle is defined by $\tan\theta=y/x$ SM. Here, the parameter $\sigma$ in the expressions (as opposed to the subscripts) takes values $\sigma=\pm 1$, corresponding to spin-up $(\uparrow)$ and spin-down $(\downarrow)$ states, respectively. The $d_{xy}$-symmetry of the altermagnetic Fermi surface is reflected by the relation $M_{x}\bm{k}_{\sigma}(\omega,\theta)M_{x}^{-1}=\bm{k}_{-\sigma}(\omega,-\theta)$, where $M_{x}$ denotes mirror reflection about the $x$-axis. For a given displacement $\bm{r}$, the propagator is dominated by $\bm{k}_{\sigma}$ state whose group velocity $\bm{v}(\bm{k}_{\sigma})=\nabla_{\bm{k}_{\sigma}}\omega(\bm{k}_{\sigma})$ is aligned with $\bm{r}$.
Due to the altermagnetic spin splitting, the propagators for opposite spins are modulated by distinct factors $\bm{k}_{\sigma}\cdot\bm{r}$ in Eq. (2). As a result, when the initial spin is oriented along the $x$ direction, the electron undergoes a distinctive spin precession [cf. Fig. 1(b)], with the wave function given by
$\psi(\bm{r};\omega)=-\frac{2\sqrt{2}\pi}{\sqrt{B^{2}-\alpha_{A}^{2}}}\begin{pmatrix}K_{0}(-i\bm{k}_{\uparrow}\cdot\bm{r})\\ K_{0}(-i\bm{k}_{\downarrow}\cdot\bm{r})\end{pmatrix}$. (3)
The feature of altermagnetic spin precession becomes evident in the far-field limit $\bm{k}_{\sigma}\cdot\bm{r}\rightarrow\infty$, where the modified Bessel function exhibits the asymptotic behavior $K_{0}(-i\bm{k}_{\sigma}\cdot\bm{r})\rightarrow\sqrt{\frac{\pi}{2\bm{k}_{\sigma}\cdot\bm{r}}}e^{i(\bm{k}_{\sigma}\cdot\bm{r}+\frac{\pi}{4})}$ arfken2011mathematical. In this limit, the spatial spin distribution can be evaluated by the wave function as $(S_{x},S_{y})=(\psi^{\dagger}\sigma_{x}\psi,\psi^{\dagger}\sigma_{y}\psi)=A(\bm{r})\left[\cos(\Delta\bm{k}\cdot\bm{r}),~-\sin(\Delta\bm{k}\cdot\bm{r})\right]$, with $\Delta\bm{k}=(\Delta k_{x},\Delta k_{y})=\bm{k}_{\uparrow}-\bm{k}_{\downarrow}$ and $A=\left[(B^{2}-\alpha_{A}^{2})\sqrt{(\bm{k}_{\uparrow}\cdot\bm{r})(\bm{k}_{\downarrow}\cdot\bm{r})}/(8\pi^{3})\right]^{-1}$ the amplitude. Owing to the $d_{xy}$-symmetry of the altermagnetic spin splitting, electrons propagating along the $\pm\theta$ directions experience effective Zeeman fields of equal magnitude but opposite sign, resulting in opposite spin precessions shown in Fig. 1(b). As a result, the spatial distribution satisfies $S_{y}(x,y)=-S_{y}(x,-y)$, which can be verified in Fig. 1(c).
In addition to the symmetric structure, the distinctive spin patterns are marked by the contours defined by $S_{y}(\bm{r})=0$, corresponding to the condition $\Delta\bm{k}\cdot\bm{r}=n\pi$, with $n$ an integer; see Fig. 1(c). Notably, for $y\gg x$, we have $\Delta k_{y}\simeq 0$ and so the condition reduces to
$x_{n}=n\mathcal{T}_{x},\ \ \mathcal{T}_{x}=\frac{\lambda_{F}}{4}\sqrt{(B/\alpha_{A})^{2}-1},$ (4)
where $\lambda_{F}=2\pi\sqrt{B/2\mu}$ is the Fermi wavelength without altermagnetism. This result shows that the spin pattern exhibits periodic oscillations along the $x$ direction, with the spatial period $\mathcal{T}_{x}$ providing a direct measure of the altermagnetic spin splitting.
In typical transport experiments, electrons are injected from electrodes of finite width $W$ as shown in Fig. 1(a), referred to as line injection. Specifically, we consider injection at $x=0$ over a finite region $y\in[-W/2,W/2]$, with spins polarized along the $x$ direction. The interference between different injection points is negligible, allowing the resulting spin textures to be interpreted as a simple superposition of the contributions from all points $\bm{r}^{\prime}=(0,y^{\prime})$ along the injection line, expressed as
$\tilde{S}_{x(y)}(\bm{r})=\int_{-W/2}^{W/2}dy^{\prime}S_{x(y)}(\bm{r}-\bm{r}^{\prime}).$ (5)
The spin distribution $\tilde{S}_{y}(\bm{r})$ for the line injection is shown in Fig. 1(d). Similar to the case of point injection, the spatial profile of $\tilde{S}_{y}(\bm{r})$ remains odd with respect to the $x$-axis. Although the overall spin pattern for line injection differs significantly from that of point injection, it is noteworthy that the spin modulation along the $x$ direction near $y\to 0^{\pm}$ resembles that at $y=\pm W/2$ for point injection, as denoted by the solid/dashed arrows in Figs. 1(c,d). To understand this, we consider the case $y>0$ without loss of generality. Since $S_{y}(\bm{r};0,y^{\prime})$ is an odd function of $y-y^{\prime}$, the integral over the interval [2yW/2,W/2][2y-W/2,W/2] in Eq. (5) vanishes. Consequently, the remaining integral is over the interval [W/2,2yW/2][-W/2,2y-W/2], which reduces to the point injection at (0,W/2)(0,-W/2) as $y\to 0^{+}$. A similar argument holds for the case $y\to 0^{-}$ as well. More importantly, although the overall spin pattern for the line injection differs significantly from that of point injection, the spatial period given in Eq. (4) remains preserved at the boundaries $y=\pm W/2$ for $x\ll W$ SM, thereby facilitating its detection through transport measurements.
Figure 2: (a) Transmission probabilities from lead 1 to the upper (blue dashed line) and lower (blue solid line) probes as a function of $x$. The red solid line represents the Hall voltage. (b) Hall voltage as a function of $x$ for different altermagnetic splitting strengths $\alpha_{A}$, with its zeros marked by green points. The gray dashed lines correspond to $\tilde{x}_{n}(\alpha_{A})$. (c,d) Same plots including the dephasing effect with $\Gamma_{v}=0.03$. The relevant parameters are $B=2$, $\mu=1$, $\alpha_{A}=0.8$, and $a=0.5$. The altermagnetic region has a size of $50\times 50$, and the hopping between the leads and the scattering region is $0.1t$.
Nonlocal spin-resolved transport.—We consider spin-resolved transport measurements in the Hall bar device as illustrated in Fig. 1(a). The setup consists of a square-shaped altermagnet connected to two longitudinal leads (leads 1 and 2), along with multiple transverse narrow leads designed to probe the spatially-resolved spin information. The transport properties are calculated based on the lattice model. Specifically, the $d_{xy}$-type altermagnet can be mapped onto a 2D square lattice as hallberg2025visualization
${H}_{\text{latt}}=$ $\epsilon\sum_{\textbf{i}}{c}_{\textbf{i}}^{\dagger}{c}_{\textbf{i}}+t\sum_{\langle\textbf{i},\textbf{j}\rangle}{c}_{\textbf{i}}^{\dagger}{c}_{\textbf{j}}+t_{\alpha}\sum_{\langle\langle\textbf{i},\textbf{j}\rangle\rangle}\beta{c}_{\textbf{i}}^{\dagger}\sigma_{z}{c}_{\textbf{j}},$ (6)
where ${c}_{\textbf{i}}=({c}_{\textbf{i}\uparrow},{c}_{\textbf{i}\downarrow})^{\text{T}}$ is the spinor fermionic operator at site i, the on-site potential is given by $\epsilon=4B/a^{2}-\mu$, the nearest-neighbour hopping (denoted by $\langle\cdot\rangle$) is $t=-B/a^{2}$, and the next-nearest-neighbour hopping (denoted by $\langle\langle\cdot\rangle\rangle$) is $t_{\alpha}=\alpha_{A}/(4a^{2})$, which accounts for the altermagnetic spin splitting. The factor $\beta=\pm$ corresponds to the hopping along the $(1,1)$ and (1,1)(1,-1) directions, respectively. The model parameters are designed such that Eq. (6) reduces to Eq. (1) in the long-wavelength limit. For all the leads, we adopt an effective model of normal electrons, given by $H_{\text{lead}}=\frac{B}{2}(k_{x}^{2}+k_{y}^{2})-\mu+M\bm{\sigma}\cdot\hat{\bm{d}}$, which is also mapped onto the square lattice. In order to probe the spatial spin information, Zeeman splitting of strength $M$ is introduced to the leads with different normalized direction vectors $\hat{\bm{d}}$. Specifically, the Zeeman field in the injection lead (lead 1) is aligned along the $x$ direction, while those in all transverse detection leads are aligned along the y-y-direction, both of which are in the ferromagnetic half-metal scenario. No Zeeman splitting is present in lead 2.
Next, we study the spin-resolved transport properties using the nonequilibrium Green’s function method. According to Landauer-Büttiker formula datta1997electronic, the currents $I_{p}$ and voltages $V_{p,q}$ in different leads are related by the transmission functions $T_{pq}$ through
$I_{p}=\frac{e^{2}}{h}\sum_{q\neq p}(T_{qp}V_{p}-T_{pq}V_{q}).$ (7)
The transmission probability from lead $q$ to $p$ is calculated by $T_{pq}=\mathrm{Tr}\left[\Gamma_{p}G^{R}\Gamma_{q}G^{A}\right]$, where the retarded and advanced Green’s functions are defined by $G^{R}=\left[G^{A}\right]^{\dagger}=\left[\omega-H_{\text{latt}}-\sum_{p}\Sigma_{p}^{R}\right]^{-1}$ and the linewidth function is given by $\Gamma_{p}=i\left(\Sigma^{R}_{p}-\Sigma^{A}_{p}\right)$. Both expressions contain the self-energies $\Sigma^{R}_{p}=\left[\Sigma^{A}_{p}\right]^{\dagger}=t_{p}^{2}G^{R}_{s,p}$ due to the coupling between lead $p$ and scattering region with strength $t_{p}$, where $G^{R}_{s,p}$ is the surface Green’s function of lead $p$ sancho1985highly. In our calculations, we consider current flowing between leads 1 and 2, with $I_{1}=-I_{2}$. The upper ($u$) and lower ($l$) leads, labeled by their positions $x$, serve as voltage probes that measure the transverse voltages $V_{u,l}(x)$ under the condition of vanishing current. The voltages at all leads, including $V_{1,2}$ and $V_{u,l}(x)$, are then solved by Eq. (7).
In Fig. 2(a), we plot the spin-resolved transmission probabilities $T_{u}(x)$ and $T_{l}(x)$ from lead 1 to the upper and lower voltage probes as a function of their position $x$. The altermagnetic spin precession induces a periodic modulation of the spin overlap between the propagating electrons and the voltage probes, resulting in spatial oscillations of the transmission probabilities. By comparing with the spin textures near $y=\pm W/2$ in Fig. 1(d), one can see that the modulation of transmission probabilities faithfully reveals the underlying spin precession. In particular, electrons reaching the upper and lower probes experience opposite spin precessions, such that their spin overlaps exhibit opposite spatial modulations. This explains the nearly $\pi$ phase shift between the oscillations of $T_{u}$ and $T_{l}$ in Figs. 2(a,c). These spatial modulations of the transmission probabilities are inherited by the Hall voltage, defined as $V_{H}(x)=V_{u}(x)-V_{l}(x)$; see Fig. 2(a). The zeros of the Hall voltage occur at positions where $T_{u}=T_{l}$, corresponding to points where the spin precession angle equals $n\pi$ for both upward and downward propagations, such that their spin overlaps with the corresponding probes are identical. Interestingly, these zero points are closely related to the solutions in Eq. (4). This correspondence is illustrated in Fig. 2(b), where the Hall voltage modulations for different altermagnetic splittings are plotted, and the zero points (green dots) are compared with $\tilde{x}_{n}=1.3x_{n}$ (dashed lines), showing quantitative agreement. Here, the factor of $1.3$ arises from the reflection at the upper and lower boundaries. Since $\mathcal{T}_{x}$ is determined by the strength $\alpha_{A}$ of altermagnetic spin splitting, our scheme provides a straightforward approach for its measurement.
Dephasing and warping effects.—The above calculations assume full phase coherence. In reality, however, dephasing and spin relaxation may occur due to inelastic scattering and interactions with magnetic impurities wu2010spin; jiang2009topological. Due to the altermagnetic coupling between spin and momentum, these two effects generally occur simultaneously. To incorporate these effects, we employ Büttiker’s approach, modeling dephasing by attaching virtual voltage probes buttiker1986role; golizadeh2007nonequilibrium; zhou2022transport. Each lattice site i in the altermagnetic region is coupled to a probe represented by the self-energy $\Sigma_{\textbf{i}}^{R}=-i\Gamma_{v}/2$, where $\Gamma_{v}$ characterizes the dephasing strength, i.e., the probability for electrons to escape into the virtual probes and lose their phase information. Mathematically, the self-energies of both real and virtual leads enter the Green’s function on an equal footing, i.e., $G^{R}=\left[\omega-H_{\text{latt}}-\sum_{p}\Sigma_{p}^{R}-\sum_{\textbf{i}}\Sigma_{\textbf{i}}^{R}\right]^{-1}$, from which the transmission matrix taking into account both real and virtual leads can be obtained. Using the same current configuration as in the case without dephasing, and imposing the condition of zero current through all voltage probes, the Hall voltages are solved by Eq. (7). In our calculations, we set $\Gamma_{v}=0.03$, which corresponds to a coherent-to-total current ratio of 41%.
The transmission probabilities $T_{u,l}$ and the Hall voltage $V_{H}$ as functions of $x$ are shown in Fig. 2(c,d). Due to the virtual voltage probes, both the direct transmissions $T_{u,l}$ and the Hall voltage $V_{H}$ decay with increasing $x$. This decay occurs because electrons absorbed and reinjected by the virtual leads lose their phase and spin information. Nevertheless, the oscillation amplitude of $V_{H}$ remains visible, and its spatial period $\tilde{x}_{n}$ is unaffected [cf. Fig. 2(d)], demonstrating the robustness of the predicted effect against dephasing. Moreover, the crystalline warping that may occur in real materials does not affect our main conclusions; see Supplemental Material SM for details.
Figure 3: Transmission probabilities and Hall voltage as a function of $\alpha_{A}$ at $x=12.5$. All other parameters are the same as in Fig. 2.
Altermagnetic spin transistor.—Altermagnets have shown great potential for spintronic applications owing to their strong spin splitting, which is an order of magnitude larger than that induced by spin–orbit coupling vsmejkal2022emerging, thereby enabling spintronic devices with higher efficiency and smaller sizes. When altermagnetic spin splitting can be controlled electrically wang2024electric, the proposed device can function as a spin transistor datta1990electronic. We have seen that, for a fixed altermagnetic spin splitting $\alpha_{A}$, the spin precession gives rise to a location-dependent Hall voltage $V_{H}$, with a period determined by $\alpha_{A}$ through Eq. (4). Consequently, tuning $\alpha_{A}$ is expected to modulate both the spin precession and therefore $V_{H}$, thereby realizing a spin transistor. In Fig. 3, we show the transmissions $T_{u,l}$ and $V_{H}$ as functions of $\alpha_{A}$, where only a single pair of Hall probes is involved. The results show that the voltage drop can be efficiently tuned, enabling clear switching with a high on/off ratio.
Physically, the external electric field is produced by dual gates ($V_{tg}$ and $V_{bg}$ in Fig. 1(a)) placed on the top and bottom of the sample, which break certain crystalline symmetries and induce altermagnetic spin splitting. Taking the candidate $\mathrm{CaMnSi}$ as an example, the electric field breaks the combined parity-time-reversal symmetry while preserving $\left[\mathcal{T}||\mathcal{C}_{4z}\right]$ symmetry wang2024electric. The altermagnetic splitting originates from the potential difference between the two nonmagnetic layers and can thus be tuned and even reversed. For opposite spin splittings, the general relations $T_{u}(\alpha_{A})=T_{l}(-\alpha_{A})$ and $V_{H}(\alpha_{A})=-V_{H}(-\alpha_{A})$ hold, as confirmed in Fig. 3.
Discussions.— It is found that the altermagnetic spin splitting induces a distinctive spin precession whose sense and rate depend strongly on the electron propagation direction, thereby establishing a clear correspondence between the spin structures in real and momentum spaces. This precession leads to a spatial modulation of the Hall voltage, whose period directly reflects the magnitude of the altermagnetic spin splitting, allowing its precise determination in the transport setup. The predicted effect and its signature are shown to be robust against both dephasing and crystalline warping. For a given material, the spatial modulation of the spin texture depends on several parameters such as the altermagnetic spin splitting, effective mass, and chemical potential. To resolve the signature with high fidelity, the width of the Hall voltage probes should be smaller than, or at most comparable to, the modulation period. Although we focus here on the $d$-wave altermagnet, similar analyses can be carried out for other symmetries vsmejkal2022emerging; vsmejkal2022beyond, where intriguing spin precession is likewise expected. As this spin precession represents an intrinsic physical property of altermagnets, it has broad implications for their spintronic applications, such as generating complex spin textures, controlling magnetization dynamics, and manipulating domain walls and skyrmions manchon2019current; obata2008current.
Acknowledgements.
We thank Gu Zhang for the helpful feedback on our manuscript. This work was supported by the National Natural Science Foundation of China (Nos. 12574051 and 12222406), the Natural Science Foundation of Jiangsu Province (Nos. BK20250008 and BK20233001), the Fundamental Research Funds for the Central Universities (Nos. KG202501 and 2024300415), and the National Key Projects for Research and Development of China (No. 2022YFA1204701).